OpenStudy (anonymous):

-9k-3≤42 (Show Work)

OpenStudy (anonymous):

Add 3 to both sides, which gives us \[-9k \le 45\] Divide both sides by -9, then we get \[k \ge 5\]

OpenStudy (anonymous):

-9k-3≤42 -9k<=42+3 -9k<=45 k>=-5

OpenStudy (anonymous):

My bad \[k \ge -5\]